### v2

@@ 2.3 Simulated acceptability
How do we relate these simulated results to the observed acceptability results described in Section 1? We propose that acceptability judgments are related to the distance between the activation peak in a neural field and the centroid of the distribution of input to the neural field. In particular, we define acceptability as the reciprocal of one plus the distance between the location of the activation peak and the centroid of the input distribution, as in Eq. 7.
Thus, acceptability ranges from to , where the size of the field. Eq. 7 indexes the degree to which an activation peak is consistent with “expected” activation under this set of inputs, operationalized as the centroid of the input distribution. An actual activation peak at a given time can deviate from the “expected” activation peak because of the influence of context and noise. The degree to which the activation peak deviates from the input centroid measures the degree to which an interpretation deviates from a canonical or expected interpretation. We focus here on deviation in the conn field, because the CA field behaves similarly across all simulations under this parameter set. In our simulation of a canonical alienable possession interpretation of have (Figure 7), all of the inputs to the conn field are centered at , the center of the field (mid conn). Thus, the centroid of the input to the conn field is . As mentioned above, under this set of inputs, the model will always tend to form an activation peak near the center of the field, leading to high acceptability. On that particular simulation, the location of the peak was , leading to acceptability .
@@
We note that, while the definition of acceptability in Eq. 7 is a heuristic, rather than a genuine neural dynamic account of acceptability judgments, the definition is generalizable. Using Eq. 7, one could calculate acceptability from any set of field inputs (corresponding to the meaning of a sentence) and any neural activation peak (corresponding to a particular interpretation of a sentence). Moreover, Eq. 7 may suggest avenues for implementing a neural dynamic model of acceptability judgments, in that the behavior of such a model should exhibit the relation defined here (i.e., higher acceptability corresponds to lower deviation between an activation peak and the input centroid). Analogous neural mechanisms may be found in the domain of movement error detection. For example, in the DIVA model of speech production (Tourville and Guenther, 2011), predicted sensory consequences of motor commands (“efference copies”: von Holst and Mittelstaedt, 1950) are compared to incoming sensory signals. A mismatch between predicted and actual sensation leads to detection of an error, and a corresponding increase in neural activation (Tourville et al., 2008). Intuitions about sentence acceptability may arise from a similar neural mechanism which compares an expected interpretation to an actual interpretation. Such a mechanism may underlie the electrophysiological difference measured between have processing in the two contexts (Zhang et al., 2018). We leave it to future work to flesh out the details of this mechanism.
@@ 2.5 Simulated response time
This account is supported by a linear regression model of simulated RT by simulated acceptability (z-scored) and context (sum-coded: possession , adjacency ). The model revealed main effects of context (, , ) and acceptability (, , ), and an interaction between context and acceptability (, , ). In order to confirm that the magnitude of the negative correlation in the adjacency context was greater than the magnitude of the positive correlation in the possession context, we multipled RT values in the possession context by , and then ran the same linear model as above. The interaction between context and acceptability remained significant (, , ), confirming a difference in the magnitudes of the correlations, independent of the directions.
@@ 3.1.2 Materials
Each experimental stimulus consisted of a pair of sentences (a context sentence followed by a target sentence) conjoined by and.888All stimuli are included in A. Every target sentence was designed to convey an adjacency interpretation of have, as in 3.1.2.
@@
| adjacency | possession |
| The maple tree has a plastic box behind it and… | The maple tree has a branch that is thick and… |
@@ 3.1.3 Procedure
The experiment began with four practice trials unrelated to the experimental stimuli. Then, each stimulus was presented twice in order to increase robustness to noise, given that the data was collected online.999An anonymous reviewer raises the potential concern of unintended priming at the second presentation of each stimulus. In order to investigate this, we also conducted all analyses separately for each presentation block. The pattern of results reported in Section 3.2 was mostly equivalent between blocks. In particular, in both blocks, we observed a participant-level correlation between AQ score and context effect on acceptability (see Section 3.2.1), and an interaction between context and acceptability in predicting reading time (see Section 3.2.2). However, there was a difference between blocks in the relationship between context, trial number, and acceptability. The effect of context on acceptability was only significant in block 2 and not block 1, and the effect of trial number on acceptability was only significant in block 1 and not block 2. In order to investigate this pattern further, we ran a single linear mixed effects model with the same structure as the model described in Section 3.2.1, except that it also included a main effect of block, as well as all interactions between block, context, and trial number. Consistent with the model reported in Table 5, we found main effects of trial number (, , ) and context=adjacency (, , ). We also found a two-way interaction between context=adjacency and trial number (, , ), a two-way interaction between block=2 and trial number (, , ), and a marginal three-way interaction between context=adjacency, block=2, and trial number (, , ). This pattern of results suggests that acceptability ratings in the possession context generally decreased over the course of block 1, before reaching a minimum in block 2. Priming would not explain this effect, since priming should lead to an increase in acceptability in block 2 relative to block 1. Therefore, we interpret the difference between blocks as primarily the result of nonlinearity in the effect of trial number. In the remainder of the paper, we report results from across the two presentation blocks. Each of the two blocks was pseudo-randomized such that no two consecutive trials were from the same condition or the same stimulus set. The presentation order of the two blocks was counter-balanced between participants. Each participant completed 160 trials in total. After completing the combined self-paced reading and acceptability judgment task, participants completed the AQ. The entire procedure was conducted in the same session in Gorilla (Anwyl-Irvine et al., 2020). The session lasted approximately 30-40 minutes.
@@ 3.1.4 Data processing
| Estimate | Std. Error | df | t value | p value | |
| char_length | |||||
| trial_num | |||||
| prec_logRT |
@@ 3.2.1 Acceptability
| Estimate | Std. Error | df | t value | p value | |
| trial number | |||||
| context=adjacency |
@@
This trend can also be seen in Figure 17, which plots mean rating (z-scored by participant) by condition against participant AQ scores, binned into quintiles, analogous to Figure 12 which displays the simulation results. While not as clean as the simulation results, there is a general trend that the difference in mean rating between conditions tends to decrease as AQ score increases.
@@ 3.2.2 Reading time
Next, we turn to the reading time results. We fit nested linear mixed effects models to the summed residualized log-transformed reading times (see Section 3.1.4) of the critical word [noun2] and each of the words in the spillover region “that is [adj]”. All models included random intercepts by item and by participant.101010Random slopes were not included because they led to model convergence issues and were also not of theoretical interest in this case. Using the same methods for fitting nested linear mixed effects models as above (Section 3.2.1), we assessed the influence of the following fixed factors on reading time: context (sum-coded: possession , adjacency ), acceptability (z-scored by subject), and their interaction. Adding a fixed factor of context did not lead to a significant improvement in fit over a baseline model with only random effects (, ). However, adding a fixed factor for acceptability rating did lead to a significant improvement in fit over the baseline model (, ). Adding a fixed factor for context back in did not lead to an improvement in fit over the model with only a fixed factor for acceptability (, ). Importantly, however, adding an interaction term did lead to a significant improvement in fit over the model with only fixed factors for main effects (, ). We report the results of the full model with both main effects and their interaction in Table 6.
| Estimate | Std. Error | df | t value | p value | |
| context | |||||
| acceptability | |||||
| context:acceptability |
@@
We fit separate linear mixed effects models to residualized log-transformed reading time for each word, with the same structure as the model reported in Table 6. There was no main effect of context on RT for the critical word [noun2] (, , ) or the following word “that” (, , ). However, at the verb “is” (two words after the critical word), there was a main effect of context on RT in the opposite direction to the prediction (, , ). Not until [adj], three words after the critical word, did we observe a main effect of context on RT in the predicted direction (, , ). We hypothesize that the unpredicted effect on “is” was due to unintended lexical priming in the possession context, since both the possession context sentence and the target sentence included a relative clause beginning with “that is”, while the adjacency context sentence did not include this phrase. This low-level similarity between the possession context sentence and the target sentence, which was not present in the stimuli of Zhang et al. (2018), likely obscured the main effect of context on RT when measured across the entire critical region. This said, our replication of the acceptability effect and the by-participant correlation between the acceptability effect and AQ score, as well as our confirmation of the predicted interaction between context and acceptability in predicting reading time, suggests that this difference in stimulus design did not otherwise confound the results of the experiment.
@@ 3.3 Discussion
The acceptability results replicate previously reported results regarding contextual facilitation of adjacency have readings. Specifically, the acceptability of sentences with have and two unrelated, inanimate nouns was improved following a context sentence that conveyed an adjacency have reading, relative to a context sentence that conveyed a possession have reading (Figure 15). Moreover, individual variation in the magnitude of this improvement was predicted by AQ scores: individuals with higher AQ scores (lower context-sensitivity) showed reduced contextual facilitation (Figures 16, 17).
With respect to reading time, the predicted main effect of context on reading time was observed on the last word in the critical region; at the previous word, the opposite effect was observed (Figure 20), arguably due to unintended priming in the possession context, since both the possession context sentence and the target sentence contained the phrase “that is”. Finally, the predicted relationship between context, acceptability, and reading time was partially supported (Table 6, Figure 18). As predicted, there was a significant main effect of acceptability on reading time, a significant interaction between context and acceptability in predicting reading time, and a significant negative correlation between reading time and acceptability in the adjacency context. However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We discuss a possible explanation for this unpredicted result in the following section.
@@ 4.1 Summary & discussion
In the model, lexical meaning is a coupling pattern between a neural node representing the lexical item and dynamic neural fields (DNFs) governing interpretation on continuous semantic dimensions. Interpretation occurs in time as activation of the lexical node causes peaks of activation in the semantic DNFs. The locations of the peaks in feature space correspond to the content of the interpretation. Dependencies between semantic dimensions, schematized in Figures 1 and 2, are modeled as coupling patterns between DNFs.
Simulations from the model captured known empirical effects. In particular, the specific reading evoked by English have was influenced by preceding context. While adjacency readings were overall more likely than possession readings (because the two nouns in the target sentence were unrelated and both inanimate), the likelihood of a possession reading was increased following a possession context, analogous to syntactic priming effects in comprehension (e.g., Arai et al., 2007; Ledoux et al., 2007; Tooley and Traxler, 2010). Moreover, adjacency readings were generally reached more quickly than possession readings, as reflected by the main effect of context on response time. We also related the simulation results to acceptability judgments through a generalizable measure (Eq. 7), deriving the empirical effect of context on acceptability (Zhang et al., 2018). The effects of the context sentence on target sentence interpretation occurred in the model because activation states persist in time, continuing to bias interpretation until they return to a resting state.111111The model parameter , which controls the rate of field evolution, modulates the temporal extent of contextual bias on interpretation.
@@
Finally, model simulations predicted an interaction between context and acceptability in predicting response time (see Figure 14). This prediction has not previously been reported or empirically tested (although other models may generate the same prediction under certain assumptions; see Section 4.2.2). Our model generated this prediction in the following way. The adjacency context sentence reinforced the overall bias for an adjacency interpretation of the target sentence, leading to a negative correlation between acceptability and response time. The possession context sentence introduced a bias towards a possession interpretation of the target sentence, competing with the overall bias towards an adjacency interpretation. This led to a positive correlation between acceptability and response time in the possession context, but of a weaker magnitude than the negative correlation observed in the adjacency context. This prediction was partially confirmed by the results of a behavioral experiment combining self-paced reading and acceptability judgments. Specifically, results showed an interaction between context and acceptability in predicting reading time on a trial by trial basis, such that the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context (Figure 18). Also as predicted, there was a significant negative main effect of acceptability on reading time, and a negative correlation between reading time and acceptability in the adjacency context.
@@
However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We attribute this discrepancy to a bias introduced by the experimental task. While our simulated acceptability measure is derived only from the interpretation of the target sentence (which is still, of course, influenced by the context sentence), participants were instructed to rate the acceptability of both sentences together. The perceived (in)congruity of the context and target sentences may have increased the magnitude of the negative correlation between acceptability and reading time across both contexts, eliminating the predicted (but already weak) positive correlation in the possession context, since this potential factor was not incorporated in our model simulations. This said, the key finding, predicted by the model, remains that context modulates the relationship between acceptability and reading time. We leave for future work the more detailed neural modeling of acceptability judgments (see Section 2.3) that this finding motivates.

### v3

@@ 2.3 Simulated acceptability
How do we relate these simulated results to the observed acceptability results described in Section 1? Modeling human acceptability judgments is far from trivial. The process of making an acceptability judgment is influenced by a variety of linguistic and non-linguistic factors (for discussion of some of these issues, see Cowart (1997); Lau et al. (2014); Schütze (1996); Sprouse et al. (2018)). Here, we propose a particular operationalization of acceptability that is intended to capture meaningful variation in the context of this complexity. We propose that acceptability judgments are related to the distance between the activation peak in a neural field and the centroid of the distribution of input to the neural field. More specifically, we define acceptability as the reciprocal of one plus the distance between the location of the activation peak and the centroid of the input distribution, as in Eq. 1.
Thus, acceptability ranges from to , where the size of the field. Eq. 1 indexes the degree to which an activation peak is consistent with “expected” activation under this set of inputs, operationalized as the centroid of the input distribution. An actual activation peak at a given time can deviate from the “expected” activation peak because of the influence of context and noise. The degree to which the activation peak deviates from the input centroid measures the degree to which an interpretation deviates from a canonical or expected interpretation.
According to this definition, a separate acceptability measure can be calculated for each field that (1) stabilizes in an “on” state (i.e., has an activation peak) and (2) receives input from the relevant lexical items (in this case, have and its two arguments). In this study, we model two fields underlying interpretation of the target sentences: one representing connectedness (conn) and one representing control asymmetry (CA). However, under our simulation parameters (which are designed to mirror interpretation of the experimental sentences, described below), the only field in which the activation peak systematically diverges from the input distribution is the conn field. The location of the activation peak in the CA field does not vary systematically across trials. For this reason, in our calculation of acceptability, we focus only on the conn field. This yields the same qualitative results as averaging across acceptability derived from each field; including the CA field would simply introduce some additional random noise into the final average acceptability value, obscuring the effects of interest.101010For other studies that use our definition of acceptability, we recommend averaging acceptability across only those fields in which the location of the activation peak is expected to vary systematically across trials. The purpose of this recommendation is to minimize noise in the comparison between simulated and measured acceptability, given that measured acceptability is necessarily influenced by a variety of noise sources which are not explicitly modeled, as mentioned above.
In our simulation of a canonical alienable possession interpretation of have (Figure 5), all of the inputs to the conn field are centered at , the center of the field (mid conn). Thus, the centroid of the input to the conn field is . As mentioned above, under this set of inputs, the model will always tend to form an activation peak near the center of the field, leading to high acceptability. On that particular simulation, the location of the peak was , leading to acceptability .
@@
We note that, while the definition of acceptability in Eq. 1 is a heuristic, rather than a genuine neural dynamic account of acceptability judgments, the definition is generalizable. Using Eq. 1, one could calculate acceptability from any set of field inputs (corresponding to the meaning of a sentence) and any (set of) neural activation peak(s) (corresponding to a particular interpretation of a sentence). Moreover, Eq. 1 may suggest avenues for implementing a neural dynamic model of acceptability judgments, in that the behavior of such a model should exhibit the relation defined here (i.e., higher acceptability corresponds to lower deviation between an activation peak and the input centroid). Analogous neural mechanisms may be found in the domain of movement error detection. For example, in the DIVA model of speech production (Tourville and Guenther, 2011), predicted sensory consequences of motor commands (“efference copies”: von Holst and Mittelstaedt, 1950) are compared to incoming sensory signals. A mismatch between predicted and actual sensation leads to detection of an error, and a corresponding increase in neural activation (Tourville et al., 2008). Intuitions about sentence acceptability may arise from a similar neural mechanism which compares an expected interpretation to an actual interpretation. Such a mechanism may underlie the electrophysiological difference measured between have processing in the two contexts (Zhang et al., 2018). We leave it to future work to flesh out the details of this mechanism.
@@ 2.5 Simulated response time
This account is supported by a linear regression model of simulated RT by simulated acceptability (z-scored) and context (sum-coded: possession , adjacency ). The model revealed main effects of context (, , ) and acceptability (, , ), and an interaction between context and acceptability (, , ). In order to confirm that the magnitude of the negative correlation in the adjacency context was greater than the magnitude of the positive correlation in the possession context, we multiplied RT values in the possession context by , and then ran the same linear model as above. The interaction between context and acceptability remained significant (, , ), confirming a difference in the magnitudes of the correlations, independent of the directions.
@@ 3.1.2 Materials
Each experimental stimulus consisted of a pair of sentences (a context sentence followed by a target sentence) conjoined by and.111111All stimuli are included in B. Every target sentence was designed to convey an adjacency interpretation of have, as in 3.1.2.
@@ 3.1.3 Procedure
The experiment began with four practice trials unrelated to the experimental stimuli. Then, each stimulus was presented twice in order to increase robustness to noise, given that the data was collected online.121212An anonymous reviewer raises the potential concern of unintended priming at the second presentation of each stimulus. In order to investigate this, we also conducted all analyses separately for each presentation block. The pattern of results reported in Section 3.2 was mostly equivalent between blocks. In particular, in both blocks, we observed a participant-level correlation between AQ score and context effect on acceptability (see Section 3.2.1), and an interaction between context and acceptability in predicting reading time (see Section 3.2.2). However, there was a difference between blocks in the relationship between context, trial number, and acceptability. The effect of context on acceptability was only significant in block 2 and not block 1, and the effect of trial number on acceptability was only significant in block 1 and not block 2. In order to investigate this pattern further, we ran a single linear mixed effects model with the same structure as the model described in Section 3.2.1, except that it also included a main effect of block, as well as all interactions between block, context, and trial number. Consistent with the model reported in Table 4, we found main effects of trial number (, , ) and context=adjacency (, , ). We also found a two-way interaction between context=adjacency and trial number (, , ), a two-way interaction between block=2 and trial number (, , ), and a marginal three-way interaction between context=adjacency, block=2, and trial number (, , ). This pattern of results suggests that acceptability ratings in the possession context generally decreased over the course of block 1, before reaching a minimum in block 2. Priming would not explain this effect, since priming should lead to an increase in acceptability in block 2 relative to block 1. Therefore, we interpret the difference between blocks as primarily the result of nonlinearity in the effect of trial number. In the remainder of the paper, we report results from across the two presentation blocks. Each of the two blocks was pseudo-randomized such that no two consecutive trials were from the same condition or the same stimulus set. The presentation order of the two blocks was counter-balanced between participants. Each participant completed 160 trials in total. After completing the combined self-paced reading and acceptability judgment task, participants completed the AQ. The entire procedure was conducted in the same session in Gorilla (Anwyl-Irvine et al., 2020). The session lasted approximately 30-40 minutes.
@@ 3.2.1 Acceptability
This trend can also be seen in Figure 15, which plots mean rating (z-scored by participant) by condition against participant AQ scores, binned into quintiles, analogous to Figure 10 which displays the simulation results. The difference in mean rating between conditions tends to decrease as AQ score increases, although this trend is not as clean as in the simulation results. We revisit this result in Section 3.3.
@@ 3.2.2 Reading time
Next, we turn to the reading time results. We fit nested linear mixed effects models to the summed residualized log-transformed reading times (see Section 3.1.4) of the critical word [noun2] and each of the words in the spillover region “that is [adj]”. All models included random intercepts by item and by participant.131313Random slopes were not included because they led to model convergence issues and were also not of theoretical interest in this case. Using the same methods for fitting nested linear mixed effects models as above (Section 3.2.1), we assessed the influence of the following fixed factors on reading time: context (sum-coded: possession , adjacency ), acceptability (z-scored by subject), and their interaction. Adding a fixed factor of context did not lead to a significant improvement in fit over a baseline model with only random effects (, ). However, adding a fixed factor for acceptability rating did lead to a significant improvement in fit over the baseline model (, ). Adding a fixed factor for context back in did not lead to an improvement in fit over the model with only a fixed factor for acceptability (, ). Importantly, however, adding an interaction term did lead to a significant improvement in fit over the model with only fixed factors for main effects (, ). We report the results of the full model with both main effects and their interaction in Table 5.
@@
We fit separate linear mixed effects models to residualized log-transformed reading time for each word, with the same structure as the model reported in Table 5. There was no main effect of context on RT for the critical word [noun2] (, , ) or the following word “that” (, , ). However, at the verb “is” (two words after the critical word), there was a main effect of context on RT in the opposite direction to the prediction (, , ). Not until [adj], three words after the critical word, did we observe a main effect of context on RT in the predicted direction (, , ). We discuss a possible explanation for this pattern in Section 3.3.
@@ 3.3 Discussion
The acceptability results replicate previously reported results regarding contextual facilitation of adjacency have readings (Figure 13), as well as individual variation in the magnitude of facilitation (Figures 14, 15). The qualitative similarity between the effect of AQ score on measured acceptability ratings (Figures 14, 15) and the effect of field coupling strength on simulated acceptability (Figure 10) provides some support for our proposal that measured AQ scores are related to field coupling strength (see Section 2.4).
With respect to reading time, the predicted main effect of context on reading time was observed on the last word in the critical region; at the previous word (“is”), the opposite effect was observed (Figure 18). We hypothesize that the unpredicted effect on “is” was due to unintended lexical priming in the possession context, since both the possession context sentence and the target sentence included a relative clause beginning with “that is”, while the adjacency context sentence did not include this phrase. This low-level similarity between the possession context sentence and the target sentence, which was not present in the stimuli of Zhang et al. (2018), likely obscured the main effect of context on RT when measured across the entire critical region. This said, our replication of the acceptability effect and the by-participant correlation between the acceptability effect and AQ score, as well as our confirmation of the predicted interaction between context and acceptability in predicting reading time, suggests that this difference in stimulus design did not otherwise confound the results of the experiment.
Finally, the predicted relationship between context, acceptability, and reading time was partially supported (Table 5, Figure 16). As predicted, there was a significant main effect of acceptability on reading time, a significant interaction between context and acceptability in predicting reading time, and a significant negative correlation between reading time and acceptability in the adjacency context. However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We discuss a possible explanation for this unpredicted result in Section 4.1.
@@ 4.1 Summary & discussion
In the model, lexical meaning is a coupling pattern between a neural node representing the lexical item and dynamic neural fields (DNFs) governing interpretation on continuous semantic dimensions. Interpretation occurs in time as activation of the lexical node causes peaks of activation in the semantic DNFs. The locations of the peaks in feature space correspond to the content of the interpretation. Dependencies between semantic dimensions, schematized in Figures 1 and 2, are modeled as coupling patterns between DNFs, shown in Figure 4.
Simulations from the model captured known empirical effects. In particular, the specific reading evoked by English have was influenced by preceding context. While adjacency readings were overall more likely than possession readings (because the two nouns in the target sentence were unrelated and both inanimate), the likelihood of a possession reading was increased following a possession context, analogous to syntactic priming effects in comprehension (e.g., Arai et al., 2007; Ledoux et al., 2007; Tooley and Traxler, 2010). Moreover, adjacency readings were generally reached more quickly than possession readings, as reflected by the main effect of context on response time. We also related the simulation results to acceptability judgments through a generalizable measure (Eq. 1), deriving the empirical effect of context on acceptability (Zhang et al., 2018). The effects of the context sentence on target sentence interpretation occurred in the model because activation states persist in time, continuing to bias interpretation until they return to a resting state.141414The model parameter , which controls the rate of field evolution, modulates the temporal extent of contextual bias on interpretation.
@@
Finally, under our proposed definition of acceptability (see Section 2.3), model simulations predicted an interaction between context and acceptability in predicting response time (see Figure 12). This prediction has not previously been reported or empirically tested (although other models may generate the same prediction under certain assumptions; see Section 4.2.2). Our model generated this prediction in the following way. Recall that, during target sentence processing, activation trajectories in the conn DNF were generally more likely to converge to an activation peak at the low end of the DNF (i.e., an adjacency interpretation), due to the weak input from the two nouns in the target sentence (see Section 2.2). In other words, independent of the influence of the context sentence, the adjacency attractor was slightly stronger than the possession attractor. For this reason, there was a main negative effect of simulated acceptability on simulated response time, such that when a more acceptable (adjacency) reading was reached, it tended to be reached more quickly. The context sentence additionally biased target sentence processing by exerting an influence on the initial neural activation state at the onset of target sentence processing. The adjacency context sentence reduced the distance in state space between the initial activation state and the adjacency attractor. This reinforced the overall bias for activation trajectories to converge to the adjacency attractor, magnifying the overall negative relationship between acceptability and response time. At the end of the possession context sentence, on the other hand, the neural activation state in the conn DNF was closer in state space to the possession attractor. This increased the likelihood of convergence to the possession attractor, competing with the overall bias towards the adjacency attractor. As a result, in the possession context, there was a positive correlation between simulated acceptability and simulated response time, such that when a more acceptable (adjacency) reading was reached, it tended to be reached more slowly. However, the magnitude of the positive correlation in the possession context was of a weaker magnitude than the negative correlation observed in the adjacency context. The reason for this is that the correlation in the adjacency context reflects the combination of two reinforcing influences, while the correlation in the possession context reflects a conflict between two competing influences.
The predicted relationship between context, acceptability, and response time was partially confirmed by the results of a behavioral experiment combining self-paced reading and acceptability judgments. Specifically, results showed an interaction between context and acceptability in predicting reading time on a trial by trial basis, such that the relationship between acceptability and reading time was more negative in the adjacency context relative to the possession context (Figure 16). Also as predicted, there was a significant negative main effect of acceptability on reading time, and a negative correlation between reading time and acceptability in the adjacency context.
@@
However, the predicted positive correlation between reading time and acceptability in the possession context was not observed. We attribute this discrepancy to a bias introduced by the experimental task. While our simulated acceptability measure is derived only from the interpretation of the target sentence (which is still, of course, influenced by the context sentence), participants were instructed to rate the acceptability of both sentences together. The perceived (in)congruity of the context and target sentences may have increased the magnitude of the negative correlation between acceptability and reading time across both contexts, eliminating the predicted (but already weak) positive correlation in the possession context, since this potential factor was not incorporated in our model simulations. This said, the key finding, predicted by the model, is that context modulates the relationship between acceptability and reading time. We leave for future work the more detailed neural modeling of acceptability judgments (see Section 2.3) that may help shed light on this issue.